OpenStudy (anonymous):

Part 1: Write the equation of the line that passes through the points (–1, –4) and (2, 5) in point-slope form. Part 2: Using complete sentences, explain whether or not it matters which point is used in the final answer. Also explain why you chose the point you did.

OpenStudy (anonymous):

@texaschic101 last one

OpenStudy (texaschic101):

first we find the slope... slope(m) = (y2 - y1) / (x2 - x1) (-1,4) x1 = -1 and y1 = 4 (2,5) x2 = 2 and y2 = 5 time to sub... m = (5 - 4) / (2 - (-1) m = 1/(2 + 1) m = 1/3 point slope form : y - y1 = m(x - x1) using points (2,5) y - 5 = 1/3(x - 2) <-- this is point slope equation y - 5 = 1/3x - 2/3 y = 1/3x - 2/3 + 5 y = 1/3x - 2/3 + 15/3 y = 1/3x + 13/3 using points (-1,4) y - y1 = m(x - x1) y - 4 = 1/3(x - (-1) <-- this is point slope equation y - 4 = 1/3x + 1/3 y = 1/3x + 1/3 + 4 y = 1/3x + 1/3 + 12/3 y = 1/3x + 13/3 It does not matter which points you use because as you can see, you end up with the same answer. I just prefer, if possible, to use the points that are positive. You may choose either points for any reason you choose.